A researcher developed a regression model to predict the tear rating of a bag of coffee based on the plate gap in​ bag-sealing equipment. Data were collected on

2828

bags in which the plate gap was varied. An analysis of variance from the regression showed that

b1equals=0.79240.7924

and

Upper S Subscript b 1Sb1equals=0.20940.2094.

a. At the 0.05 level of​ significance, is there evidence of a linear relationship between the plate gap of the​ bag-sealing machine and the tear rating of a bag of​ coffee?

b. Construct a​ 95% confidence interval estimate of the population​ slope,

betaβ1.
Answer #1

Answer:

a)

Given,

b1 = 0.7924

Sb1 = 0.2094

Now,

t = b1/Sb1

substitute the known values

t = 0.7924/.2094

test statistic = 3.784

Degree of freedom = n - 2

= 28 - 2

= 26

So the critical value is tcrit = 2.056

Here t > tcrit , so there is a linear relationship b/w plate of the​ bag-sealing machine and the tear rating of a bag of​ coffee.

b)

To determine the 95% confidence interval

Now CI = b1 +/- t0.05,26 Sb1

substitute the known values

= 0.7924 +/- 2.056*0.2094

= 0.7924 +/- 0.4305264

CI = (0.3619 , 1.2229)